OpenStudy (anonymous):

two plains leave the same airport and go in opposite directions at the same time, at 600 mph and 350 respectively. in how many hours will they be 2850 miles apart?

OpenStudy (anonymous):

600x+350x=2850

OpenStudy (anonymous):

try planes not plains

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

can you still do the problem?

OpenStudy (anonymous):

600x+350x = 2850 miles 950x = 2850 miles x = 2850/950 x = 3 They will be 2850 miles apart at exactly 3 hours

OpenStudy (anonymous):

Trying solving for x

OpenStudy (anonymous):

ok thanks dude!